A lab technician uses a 0.25-mol/L sodium carbonate stock solution, Na2CO3(aq), to prepare 250 mL of 0.010-mol/L solution. How much of the stock solution should be used?


1
Expert's answer
2021-03-22T06:38:17-0400

The dilution coefficient equals:

n = c1/c2 = 0.25 mol/L / 0.010 mol/L = 25

From here, the volume of the stock solution that must be used is:

V1 = V2 / n = 10 mL


Answer: 10 mL